OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. a15 = -53, a16 = -5

hartnn (hartnn):

can you find common difference 'd' from that information ?

OpenStudy (anonymous):

I don't know how to find it?

OpenStudy (anonymous):

These are my options for an answer: a. an = -725 - 48(n - 1) b. an = -725 + 48(n + 1) c. an = -725 + 48(n - 1) d. an = -725 - 48(n + 1)

OpenStudy (anonymous):

How do they get the number -725 and -48?

hartnn (hartnn):

the common difference 'd' is the difference between consecutive terms! so, \(\Large a_{16}- a_{15} = d\) first find d

OpenStudy (anonymous):

Oh! d = 48

hartnn (hartnn):

yes!

OpenStudy (anonymous):

a is correct

hartnn (hartnn):

now u just need 1st term since we know 15th term, we use the formula \(\Large a_{15} = a_1 + (15-1)\times 48\)

OpenStudy (anonymous):

Okay, thanks. @HARSH123 @hartnn What is a1? How do I do that part?

hartnn (hartnn):

a_1 is the first term

hartnn (hartnn):

we already know a15 as -53 so find a_1 from \(-53 = a_1 +(15-1)\times 48\)

OpenStudy (anonymous):

\[-53 = a _{1} + 14 * 48 \]

OpenStudy (anonymous):

\[-53 = a _{1} + 672\]

hartnn (hartnn):

yes, so a1= ... ?

OpenStudy (anonymous):

Um. Do I say 672- (-53) = a1?

OpenStudy (anonymous):

Oh yes. But that gives me +725

hartnn (hartnn):

-53 = a1 +672 a1 = -52 -672 = -725

OpenStudy (anonymous):

I'm confused, how do I know to make a1's equation negative?

hartnn (hartnn):

\(-53 = a _{1} + 672 \\ \text{subtracting 672 on both sides} \\ -53 -672 = a_1 \) a1 was always negative in this case

OpenStudy (anonymous):

Ohhhh I see! Thank you so much! So the correct answer is: an = -725 - 48(n - 1)

hartnn (hartnn):

the formula is \(a_n = a_1 +(n-1)d \\ so, a_n = -725 \large + 48(n-1)\)

hartnn (hartnn):

option c

OpenStudy (anonymous):

You're awesome, thank you :)

hartnn (hartnn):

and you too are good! welcome ^_^